1. how to make a high level analysis of a company's balance sheet and income statement using vertical and horizontal analysis
1
Expert's answer
2020-12-17T07:10:00-0500

Solution:

Horizontal analysis is a trend analysis used in financial statement analysis to compare historical data, such as line items over a number of accounting periods. Under the income statement, you can compare each line item's current year figure and prior-year figure. For example, you can compare current year sales and previous year sales figures, including other line items in the income statement. The same applies to the balance sheet statement, where you can compare each line item's current figures and previous figures. For example, you can compare current inventory figures with previous year figures, including other line items in the balance sheet.

Horizontal analysis can either use absolute comparisons or percentage comparisons, which are widely utilized. Percentage comparisons are where the numbers in each succeeding period are expressed as a percentage of the amount in the baseline year. The formula is as follows:


"Horizontal\; analysis\; formula\;(\%) =\frac{(Amount\; in\; comparison\; year - Amount \;in \;base \;year) }{Amount \;in\; base \;year} \times 100"



Vertical analysis is a method of financial statement analysis in which a specific line item is listed as a percentage of a base figure within a particular statement. In the income statement, each line item is stated as a percentage of sales, while line items on a balance sheet statement are stated as a percentage of total assets or liabilities and shareholder’s equity.

The vertical analysis uses percentage comparisons. The formula is as follows for both income statement and balance sheet statement:



"Vertical\; analysis\; formula\;(Income\; statement) =\frac{Income \;statement\; line\; item}{Total\; sales} \times 100"


"Vertical\; analysis\; formula\;(Balance \; sheet\;statement) =\frac{Balance\;sheet \;statement\; line\; item}{Total\; Assets(Liabilities) } \times 100"